# thanks: Corresponding author

### v1

@@ I Introduction
Problems with applying noniterative corrections to CCSD in situations where $T_{n}$ components with $n>2$, such as $T_{3}$, are not only large and nonperturbative, but also strongly coupled to their lower-rank $T_{1}$ and $T_{2}$ counterparts, have motivated us to develop the generalization of the biorthogonal moment expansions, which in the past resulted in the CR-CC approaches, such as CR-CC(2,3) and its excited-state and higher-order extensions, Piecuch and Włoch (2005); Piecuch et al. (2006); Włoch et al. (2006, 2007); Piecuch et al. (2006, 2009); Fradelos et al. (2011); Piecuch et al. (2007, 2008); Horoi et al. (2007); Bauman et al. (2017); Magoulas et al. (2018) to unconventional truncations in the cluster and EOMCCEmrich (1981, 1981); Geertsen et al. (1989); Stanton and Bartlett (1993) excitation operators, designated as CC($P$;$Q$). Shen and Piecuch (2012, 2012, 2012); Bauman et al. (2017); Magoulas et al. (2018); Yuwono et al. (2019); Deustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021); Chakraborty et al. (2022); Gururangan et al. (2021); Gururangan and Piecuch (2023) By incorporating the dominant contributions to the higher–than–two-body clusters into the iterative steps, so that $T_{1}$ and $T_{2}$ amplitudes can relax compared to their CCSD values when $T_{n}$ components with $n>2$ become more substantial, and correcting the results for the remaining many-electron correlation effects of interest using suitably defined moment expansions, the CC($P$;$Q$) formalism provides us with the opportunity to converge or accurately approximate the parent high-level CCSDT, CCSDTQ, and similar energetics at small fractions of the computational costs, even when noniterative corrections to CCSD fail or struggle. Focusing on full CCSDT, which provides the parent data for the lowest singlet and triplet PECs of cyclobutadiene examined in this work, a few different CC($P$;$Q$) approaches designed to converge CCSDT energetics have been developed. In the initial, active-orbital-based, variant of CC($P$;$Q$), abbreviated as CC(t;3), which is part of the larger CC(t;3), CC(t,q;3), CC(t,q;3,4), CC(q;4), etc. hierarchy,Shen and Piecuch (2012, 2012, 2012); Bauman et al. (2017); Magoulas et al. (2018); Yuwono et al. (2019) the leading $T_{3}$ amplitudes that enter the iterative steps preceding the determination of the CC($P$;$Q$) corrections are obtained using the active-space CCSDt approach.Oliphant and Adamowicz (1992, 1993); Piecuch et al. (1993); Piecuch and Adamowicz (1994, 1995); Ghose et al. (1995); Adamowicz et al. (1998); Piecuch et al. (1999, 1999); Piecuch (2010) In the more black-box semi-stochastic Deustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021); Chakraborty et al. (2022) and CIPSI-driven CC($P$;$Q$)Gururangan et al. (2021) approaches aimed at converging CCSDT, the dominant triply excited cluster amplitudes are identified with the help of CI Booth et al. (2009); Cleland et al. (2010); Dobrautz et al. (2019); Ghanem et al. (2019, 2020) or CCThom (2010); Franklin et al. (2016); Spencer and Thom (2016); Scott and Thom (2017) Quantum Monte Carlo wave function propagations in the many-electron Hilbert space, in the former case, and the sequences of Hamiltonian diagonalizations constructed in the CIPSI algorithmHuron et al. (1973); Garniron et al. (2017, 2019) in the case of the latter method. In the recently introduced adaptive CC($P$;$Q$) formalism, Gururangan and Piecuch (2023) the triply excited determinants and amplitudes defining the leading $T_{3}$ contributions in the iterative steps of CC($P$;$Q$) calculations are identified using the intrinsic structure of the CC($P$;$Q$) energy corrections.
@@
In this study, we focus on the CIPSI-driven CC($P$;$Q$) methodology of Ref. Gururangan et al., 2021. Our main goal is to answer the question how efficient this methodology is in converging the lowest singlet and triplet potentials of cyclobutadiene along the $D_{\rm 2h}$-symmetric automerization coordinate resulting from the highly accurate CCSDT computations shown in Figure 1. As demonstrated in our initial study,Gururangan et al. (2021) the CIPSI-driven CC($P$;$Q$) approach is capable of producing the near-CCSDT energetics for singlet electronic states using tiny fractions of triply excited cluster amplitudes in the iterative parts of the CC($P$;$Q$) algorithm that are considerably smaller than those used in the analogous semi-stochastic and active-orbital-based CC($P$;$Q$) considerations. One of the key objectives of this work is to determine if similar observations apply to the CIPSI-driven CC($P$;$Q$) calculations for the lowest singlet and triplet potentials of cyclobutadiene along its automerization coordinate and the gap between them. The role of the CC($P$;$Q$) moment corrections in improving the resulting energetics is discussed as well.
@@ II Theory and Computational Details
One of the main advantages of the CC($P$;$Q$) methodology is its flexibility. In particular, we can make a wide variety of conventional as well as unconventional choices of the $P$ and $Q$ spaces, adjusting them to the nature of the electronic states of interest and adopting different numerical procedures in their construction. Conventional choices for the $P$ and $Q$ spaces, based on the many-body ranks of the determinants included in them, result in the left-eigenstate CR-CC methods, such as the CR-CC(2,3) approach discussed in the Introduction, in which the former space consists of all singly and doubly excited determinants and the latter space is spanned by all triples. We can, however, also make unconventional choices, including those adopted in the CC(t;3), CC(t,q;3), CC(t,q;3,4), CC(q;4), etc. hierarchy Shen and Piecuch (2012, 2012, 2012); Bauman et al. (2017); Magoulas et al. (2018); Yuwono et al. (2019) and the semi-stochastic, Deustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021); Chakraborty et al. (2022) adaptive,Gururangan and Piecuch (2023) and CIPSI-drivenGururangan et al. (2021) CC($P$;$Q$) methods, mentioned in the Introduction as well, in which the suitably chosen subsets of higher–than–doubly excited determinants are incorporated into the underlying $P$ spaces, in addition to all singles and doubles, to relax the lower-rank $T_{1}$ and $T_{2}$ clusters in the presence of their higher-rank counterparts, such as the leading $T_{3}$ contributions, which the CCSD(T), CR-CC(2,3), $\Lambda$-CCSD(T), and similar approaches are not designed to do. Having some higher–than–doubly excited determinants in the $P$ space provides us with a straightforward and computationally efficient mechanism to account for the coupling between the lower- and higher-order components of the cluster operator, which cannot be neglected when $T_{n}$ contributions with $n>2$, such as $T_{3}$, become large and nonperturbative, as is the case when the automerization barrier region of the lowest-energy singlet potential of cyclobutadiene is examined. This, in turn, allows us to recover the full CCSDT, CCSDTQ, and similar energetics without running into the very expensive, often prohibitive, computational costs associated with the high-level CC methods of this type.
@@
To produce the final wave function $|\Psi^{(\text{CIPSI})}\rangle$, needed to construct the list of higher–than–doubly excited determinants to be included in the $P$ space of a given CIPSI-driven CC($P$;$Q$) calculation, and determine the associated variational ($E_{\text{var}}$) and perturbatively corrected [$E_{\text{var}}+\Delta E^{(2)}$ or $E_{\text{var}}+\Delta E_{\text{r}}^{(2)}$] CIPSI energies, the sequence of Hamiltonian diagonalizations defining the underlying CIPSI run must be terminated. This could be done by stopping at the first iteration $k$ for which the absolute value of the second-order MBPT correction $\Delta E_{k}^{(2)}$ falls below a user-defined threshold $\eta$, but, given our interest in examining the convergence of the CIPSI-based CC($P$;$Q$) calculations toward the desired high-level CC energetics, represented in this study by CCSDT, using systematically grown $P$ spaces obtained with the help of CIPSI, in this work we follow Ref. Gururangan et al., 2021 and stop when the number of determinants in the diagonalization space equalizes or exceeds the user-defined parameter $N_{\text{det(in)}}$. To ensure that the CIPSI sequences preceding our CC($P$;$Q$) calculations did not terminate too soon, before the dimensions of terminal diagonalization spaces became greater than or equal to $N_{\text{det(in)}}$, we set the aforementioned parameter $\eta$ to 1 microhartree. As a result [putting aside the convergence threshold used in the CC($P$) iterations, which we set to $10^{-8}$ hartree], all CIPSI-driven CC($P$;$Q$) computations reported in this article, along with the underlying $P$ spaces, were controlled by a single input variable $N_{\text{det(in)}}$. In addition to $N_{\text{det(in)}}$, in presenting our CC($P$) and CC($P$;$Q$) results for the lowest singlet and triplet potentials of cyclobutadiene, we also provide information about the numbers of determinants included in the terminal CIPSI wave functions $|\Psi^{(\text{CIPSI})}\rangle$ obtained for various values of $N_{\text{det(in)}}$, designated as $N_{\text{det(out)}}$. Given our choice of the subspace enlargement parameter $f$, the $N_{\text{det(out)}}$ values characterizing the $|\Psi^{(\text{CIPSI})}\rangle$ states used to identify the triply excited determinants for inclusion in the $P$ spaces employed in our CIPSI-driven CC($P$;$Q$) computations were always between $N_{\text{det(in)}}$ and $2N_{\text{det(in)}}$. With all of this in mind, the algorithm used in the CIPSI-enabled CC($P$;$Q$) calculations reported in this study, aimed at converging the CCSDT energetics, can be summarized as follows: Gururangan et al. (2021)
@@
In order to perform the CIPSI-driven CC($P$) and CC($P$;$Q$) calculations for the singlet and triplet PECs of cyclobutadiene along the automerization coordinate investigated in this work and examine their convergence toward the CCSDT potentials shown in Figure 1, we used the computer programs described in Ref. Gururangan et al., 2021, interfaced with the RHF, ROHF, and integral transformation routines in GAMESS,Schmidt et al. (1993); Barca et al. (2020); Zahariev et al. (2023) which take advantage of our highly efficient, automatically generated, Fortran CC codes that were previously exploited in implementing the active-orbital-basedShen and Piecuch (2012, 2012, 2012); Bauman et al. (2017); Magoulas et al. (2018) and semi-stochasticDeustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021) CC($P$;$Q$) approaches. We also used our group’s computer-generated CC codes interfaced with GAMESS to produce the parent CCSDT data. The results of the DEA-EOMCC(4p-2h)$\{N_{\text{u}}\}$ calculations discussed in the Introduction, reported in Figure 1 as well, were carried out with the highly efficient DEA-EOMCC routines developed in Ref. Shen and Piecuch, 2021, which became part of the official GAMESS distribution last year.
As already alluded to above, the lists of triply excited determinants that were needed to construct the $P$ spaces for the CC($P$) and CC($P$;$Q$) calculations corresponding to the various values of the input parameter $N_{\text{det(in)}}$ were extracted from the terminal CIPSI wave functions $|\Psi^{(\text{CIPSI})}\rangle$ obtained with Quantum Package 2.0, whereas the complementary $Q$ spaces, used to determine the $\delta(\mbox{$P$;$Q$})$ corrections, consisted of the remaining triples absent in the $|\Psi^{(\text{CIPSI})}\rangle$ states. We also used Quantum Package 2.0 to obtain the $E_{\text{var}}$, $E_{\text{var}}+\Delta E^{(2)}$, and $E_{\text{var}}+\Delta E_{\text{r}}^{(2)}$ energies associated with the CIPSI runs that provided the lists of triples for the CC($P$) iterations. Our CIPSI-driven CC($P$) and CC($P$;$Q$) methods targeting CCSDT that rely on the triples lists extracted from the CIPSI calculations performed with Quantum Package 2.0 have recently been incorporated in our open-source CCpy package available on GitHub,CCp implemented using a hybrid Python-Fortran programming approach.
@@
To obtain the desired insights into the convergence of the singlet and triplet potentials of cyclobutadiene toward their CCSDT counterparts, we adopted the strategy used in Ref. Gururangan et al., 2021. Thus, for each nuclear geometry along the automerization pathway considered in this work, we carried out a series of CIPSI-driven CC($P$) and CC($P$;$Q$) calculations using the $P$ and $Q$ spaces derived from the increasingly large CIPSI wave functions obtained by varying $N_{\text{det(in)}}$ in an approximately semi-logarithmic manner. We started with $N_{\text{det(in)}}=1$, where the CIPSI wave functions $|\Psi^{(\text{CIPSI})}\rangle$ are the single determinants defining the reference states $|\Phi\rangle$ used in our CC computations (RHF in the case of the singlet and ROHF in the triplet case) and the resulting CC($P$) and CC($P$;$Q$) energies become identical to those obtained with CCSD and CR-CC(2,3), respectively, and went all the way to $N_{\text{det(in)}}=10,000,000$, to reflect on the fact that as the input variable $N_{\text{det(in)}}$ becomes increasingly large and the CIPSI wave functions capture more and more triply excited determinants, the CC($P$;$Q$) energies $E^{(P+Q)}$ approach their CCSDT parents (becoming identical to them when all triply excited determinants are captured by the $|\Psi^{(\text{CIPSI})}\rangle$ states). The energies obtained in the CC($P$) computations approach their CCSDT counterparts as well, but, as discussed in the next section, their convergence toward CCSDT is much slower than that observed in the CC($P$;$Q$) runs. The ability of the CIPSI-driven CC($P$;$Q$) methodology to generate the CCSDT-level energetics using tiny fractions of triply excited determinants in the underlying $P$ spaces captured by the relatively small Hamiltonian diagonalizations, observed in the calculations for the lowest ${}^{1}A_{g}(D_{2\text{h}})$ and ${}^{3}B_{1g}(D_{2\text{h}})$ potentials of cyclobutadiene reported in this study, results in enormous savings in the computational effort compared to CCSDT. Indeed, as explained in Ref. Gururangan et al., 2021, the CIPSI runs using smaller $N_{\text{det(in)}}$ values are much faster than those needed to reach convergence, the CC($P$) calculations using tiny fractions of triples in the $P$ space are orders of magnitude faster than the corresponding CCSDT computations (cf. Ref. Gururangan and Piecuch, 2023 for the examples of timings), and the effort involved in obtaining the noniterative $\delta(\mbox{$P$;$Q$})$ corrections is similar to the determination of the triples corrections of CR-CC(2,3) or CCSD(T).

### v2

@@ I Introduction
Problems with applying noniterative corrections to CCSD in situations where $T_{n}$ components with $n>2$, such as $T_{3}$, are not only large and nonperturbative, but also strongly coupled to their lower-rank $T_{1}$ and $T_{2}$ counterparts, have motivated us to develop the generalization of the biorthogonal moment expansions, which in the past resulted in the CR-CC approaches, such as CR-CC(2,3) and its excited-state and higher-order extensions, Piecuch and Włoch (2005); Piecuch et al. (2006); Włoch et al. (2006, 2007); Piecuch et al. (2006, 2009); Fradelos et al. (2011); Piecuch et al. (2007, 2008); Horoi et al. (2007); Bauman et al. (2017); Magoulas et al. (2018) to unconventional truncations in the cluster and EOMCCEmrich (1981, 1981); Geertsen et al. (1989); Stanton and Bartlett (1993) excitation operators, designated as CC($P$;$Q$). Shen and Piecuch (2012, 2012, 2012); Bauman et al. (2017); Magoulas et al. (2018); Yuwono et al. (2019); Deustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021); Chakraborty et al. (2022); Gururangan et al. (2021); Gururangan and Piecuch (2023); Gururangan et al. (2025) By incorporating the dominant contributions to the higher–than–two-body clusters into the iterative steps, so that $T_{1}$ and $T_{2}$ amplitudes can relax compared to their CCSD values when $T_{n}$ components with $n>2$ become more substantial, and correcting the results for the remaining many-electron correlation effects of interest using suitably defined moment expansions, the CC($P$;$Q$) formalism provides us with the opportunity to converge or accurately approximate the parent high-level CCSDT, CCSDTQ, and similar energetics at small fractions of the computational costs, even when noniterative corrections to CCSD fail or struggle. Focusing on full CCSDT, which provides the parent data for the lowest singlet and triplet PECs of cyclobutadiene examined in this work, a few different CC($P$;$Q$) approaches designed to converge CCSDT energetics have been developed. In the initial, active-orbital-based, variant of CC($P$;$Q$), abbreviated as CC(t;3), which is part of the larger CC(t;3), CC(t,q;3), CC(t,q;3,4), CC(q;4), etc. hierarchy,Shen and Piecuch (2012, 2012, 2012); Bauman et al. (2017); Magoulas et al. (2018); Yuwono et al. (2019); Gururangan et al. (2025) the leading $T_{3}$ amplitudes that enter the iterative steps preceding the determination of the CC($P$;$Q$) corrections are obtained using the active-space CCSDt approach.Oliphant and Adamowicz (1992, 1993); Piecuch et al. (1993); Piecuch and Adamowicz (1994, 1995); Ghose et al. (1995); Adamowicz et al. (1998); Piecuch et al. (1999, 1999); Piecuch (2010) In the more black-box semi-stochastic Deustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021); Chakraborty et al. (2022) and CIPSI-driven CC($P$;$Q$)Gururangan et al. (2021) approaches aimed at converging CCSDT, the dominant triply excited cluster amplitudes are identified with the help of CI Booth et al. (2009); Cleland et al. (2010); Dobrautz et al. (2019); Ghanem et al. (2019, 2020) or CCThom (2010); Franklin et al. (2016); Spencer and Thom (2016); Scott and Thom (2017) Quantum Monte Carlo wave function propagations in the many-electron Hilbert space, in the former case, and the sequences of Hamiltonian diagonalizations constructed in the CIPSI algorithmHuron et al. (1973); Garniron et al. (2017, 2019) in the case of the latter method. In the recently introduced adaptive CC($P$;$Q$) formalism, Gururangan and Piecuch (2023); Gururangan et al. (2025) the triply excited determinants and amplitudes defining the leading $T_{3}$ contributions in the iterative steps of CC($P$;$Q$) calculations are identified using the intrinsic structure of the CC($P$;$Q$) energy corrections.
@@
In this study, we focus on the CIPSI-driven CC($P$;$Q$) methodology of Ref. Gururangan et al., 2021. Our main goal is to answer the question how efficient this methodology is in converging the lowest singlet and triplet potentials of cyclobutadiene along the $D_{\rm 2h}$-symmetric automerization coordinate resulting from the high-level CCSDT computations shown in Figure 1. The ability of the CC($P$;$Q$) framework using the CIPSI algorithm to identify the leading triply excited determinants for the iterative steps of the CC($P$;$Q$) procedure to accurately approximate the CCSDT energies of the $^{1}A_{g}(D_{2\text{h}})$ and $^{3}B_{1g}(D_{2\text{h}})$ states of cyclobutadiene and the gap between, especially at the most challenging TS geometry, in calculations using basis sets larger than cc-pVDZ (represented in this work by cc-pVTZDunning (1989)) is examined as well. As shown in our initial study,Gururangan et al. (2021) the CIPSI-driven CC($P$;$Q$) approach is capable of producing the near-CCSDT energetics for singlet electronic states using tiny fractions of triply excited cluster amplitudes in the iterative parts of the CC($P$;$Q$) algorithm that are smaller than those used in the analogous semi-stochastic and active-orbital-based CC($P$;$Q$) considerations. One of the objectives of this work is to determine if similar observations apply to the CIPSI-driven CC($P$;$Q$) calculations for the lowest singlet and triplet potentials of cyclobutadiene along its automerization coordinate and the gap between them. The role of the CC($P$;$Q$) moment corrections in accelerating convergence toward CCSDT and the key elements of our improved implementation of the CIPSI-based CC($P$;$Q$) method, capable of efficiently handling small but generally spotty subsets of triply excited determinants in the underlying CC iterations, illustrated by computational timings, are discussed too.
@@ II Theory and Computational Details
One of the main advantages of the CC($P$;$Q$) methodology is its flexibility. In particular, we can make a wide variety of conventional as well as unconventional choices of the $P$ and $Q$ spaces, adjusting them to the nature of the electronic states of interest and adopting different numerical procedures in their construction. Conventional choices for the $P$ and $Q$ spaces, based on the many-body ranks of the determinants included in them, result in the left-eigenstate CR-CC methods, such as the CR-CC(2,3) approach discussed in the Introduction, in which the former space consists of all singly and doubly excited determinants and the latter space is spanned by all triples. We can, however, also make unconventional choices, including those adopted in the CC(t;3), CC(t,q;3), CC(t,q;3,4), CC(q;4), etc. hierarchy Shen and Piecuch (2012, 2012, 2012); Bauman et al. (2017); Magoulas et al. (2018); Yuwono et al. (2019); Gururangan et al. (2025) and the semi-stochastic, Deustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021); Chakraborty et al. (2022) adaptive,Gururangan and Piecuch (2023); Gururangan et al. (2025) and CIPSI-drivenGururangan et al. (2021) CC($P$;$Q$) methods, mentioned in the Introduction as well, in which the suitably chosen subsets of higher–than–doubly excited determinants are incorporated into the underlying $P$ spaces, in addition to all singles and doubles, to relax the lower-rank $T_{1}$ and $T_{2}$ clusters in the presence of their higher-rank counterparts, such as the leading $T_{3}$ contributions, which the CCSD(T), CR-CC(2,3), $\Lambda$-CCSD(T), and similar approaches are not designed to do. Having some higher–than–doubly excited determinants in the $P$ space provides us with a straightforward and computationally efficient mechanism to account for the coupling between the lower- and higher-order components of the cluster operator, which cannot be neglected when $T_{n}$ contributions with $n>2$, such as $T_{3}$, become large and nonperturbative, as is the case when the automerization barrier region of the lowest-energy singlet potential of cyclobutadiene is examined. This, in turn, allows us to recover the full CCSDT, CCSDTQ, and similar energetics without running into the very expensive, often prohibitive, computational costs associated with the high-level CC methods of this type.
@@
To produce the final wave function $|\Psi^{(\text{CIPSI})}\rangle$, needed to construct the list of higher–than–doubly excited determinants to be included in the $P$ space of a given CIPSI-driven CC($P$;$Q$) calculation, and determine the associated variational ($E_{\text{var}}$) and perturbatively corrected $E_{\text{var}}+\Delta E^{(2)}$ or $E_{\text{var}}+\Delta E_{\text{r}}^{(2)}$] CIPSI energies, the sequence of Hamiltonian diagonalizations defining the underlying CIPSI run must be terminated. This could be done by stopping at the first iteration $k$ for which the absolute value of the second-order MBPT correction $\Delta E_{k}^{(2)}$ falls below a user-defined threshold $\eta$, but, given our interest in examining the convergence of the CIPSI-based CC($P$;$Q$) calculations toward the desired high-level CC energetics, represented in this study by CCSDT, using systematically grown $P$ spaces obtained with the help of CIPSI, in this work we follow Ref. Gururangan et al., 2021 and stop when the number of determinants in the diagonalization space equalizes or exceeds the user-defined parameter $N_{\text{det(in)}}$. To ensure that the CIPSI sequences preceding our CC($P$;$Q$) calculations did not terminate too soon, before the dimensions of terminal diagonalization spaces became greater than or equal to $N_{\text{det(in)}}$, we set the aforementioned parameter $\eta$ to 1 microhartree. As a result [putting aside the convergence threshold used in the CC($P$) iterations, which we set to $10^{-7}$ hartree], all CIPSI-driven CC($P$;$Q$) computations reported in this article, along with the underlying $P$ spaces, were controlled by a single input variable $N_{\text{det(in)}}$. In addition to $N_{\text{det(in)}}$, in presenting our CC($P$) and CC($P$;$Q$) results for the lowest singlet and triplet potentials of cyclobutadiene, we also provide information about the numbers of determinants included in the terminal CIPSI wave functions $|\Psi^{(\text{CIPSI})}\rangle$ obtained for various values of $N_{\text{det(in)}}$, designated as $N_{\text{det(out)}}$. Given our choice of the subspace enlargement parameter $f$, the $N_{\text{det(out)}}$ values characterizing the $|\Psi^{(\text{CIPSI})}\rangle$ states used to identify the triply excited determinants for inclusion in the $P$ spaces employed in our CIPSI-driven CC($P$;$Q$) computations were always between $N_{\text{det(in)}}$ and $2N_{\text{det(in)}}$. With all of this in mind, the algorithm used in the CIPSI-enabled CC($P$;$Q$) calculations reported in this study, aimed at converging the CCSDT energetics, can be summarized as follows: Gururangan et al. (2021)
@@
[…]
@@
To obtain the desired insights into the convergence of the singlet and triplet potentials of cyclobutadiene toward their CCSDT counterparts, we adopted the strategy used in Ref. Gururangan et al., 2021. Thus, for each nuclear geometry along the automerization pathway considered in this work, we carried out a series of CIPSI-driven CC($P$) and CC($P$;$Q$) calculations using the $P$ and $Q$ spaces derived from the increasingly large CIPSI wave functions obtained by varying $N_{\text{det(in)}}$ in an approximately semi-logarithmic manner. We started with $N_{\text{det(in)}}=1$, where the CIPSI wave functions $|\Psi^{(\text{CIPSI})}\rangle$ are the single determinants defining the reference states $|\Phi\rangle$ used in our CC computations (RHF in the case of the singlet and ROHF in the triplet case) and the resulting CC($P$) and CC($P$;$Q$) energies become identical to those obtained with CCSD and CR-CC(2,3), respectively, and went all the way to $N_{\text{det(in)}}=10,000,000$, to reflect on the fact that as the input variable $N_{\text{det(in)}}$ becomes increasingly large and the CIPSI wave functions capture more and more triply excited determinants, the CC($P$;$Q$) energies $E^{(P+Q)}$ approach their CCSDT parents (becoming identical to them when all triply excited determinants are captured by the $|\Psi^{(\text{CIPSI})}\rangle$ states). The energies obtained in the CC($P$) computations approach their CCSDT counterparts as well, but, as discussed in the next section, their convergence toward CCSDT is much slower than that observed in the CC($P$;$Q$) runs.
The ability of the CIPSI-driven CC($P$;$Q$) methodology to generate the CCSDT-level energetics using tiny fractions of triply excited determinants in the underlying $P$ spaces captured by the relatively small Hamiltonian diagonalizations, observed in the calculations for the lowest $^{1}A_{g}(D_{2\text{h}})$ and $^{3}B_{1g}(D_{2\text{h}})$ potentials of cyclobutadiene reported in this study, results in enormous savings in the computational effort compared to CCSDT. Indeed, as explained in Ref. Gururangan et al., 2021, the CIPSI runs using smaller $N_{\text{det(in)}}$ values are much faster than those needed to reach convergence, the CC($P$) calculations using tiny fractions of triples in the $P$ space are one or more orders of magnitude faster than the corresponding CCSDT computations, and the effort involved in obtaining the noniterative $\delta(\mbox{$P$;$Q$})$ corrections is similar to the determination of the triples corrections of CR-CC(2,3) or CCSD(T). As pointed out in Refs. Yuwono et al., 2020; Deustua et al., 2021, the key to obtaining the desired computational efficiency in the CC($P$;$Q$) calculations lies in the development of an algorithm capable of offering significant speedups compared to the parent CC approach, such as CCSDT, when the lists of higher–than–doubly excited determinants included in the CC($P$) iterations do not necessarily form continuous manifolds, as is the case when these lists are created by the sequences of Hamiltonian diagonalizations of CIPSI adopted in this study and Ref. Gururangan et al., 2021, the previously employed CIQMC/CCMC propagations, Deustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021, 2021); Chakraborty et al. (2022) or the moment expansions defining the $\delta(\mbox{$P$;$Q$})$ corrections.Gururangan and Piecuch (2023); Gururangan et al. (2025) In all these cases, conventional diagrammatic (or algebraic) techniques assuming continuous excitation manifolds labeled by occupied and unoccupied orbitals from the respective ranges of indices, used to implement the standard CC methods employing rank-based truncations, no longer apply. A different programming approach is needed. The key elements of our algorithm used to implement the CC($P$) equations in which the $P$ space consists of all singly and doubly excited determinants and a generally spotty subset of triply excited determinants identified in this work by CIPSI, along with illustrative computational timings obtained using our CCpy codes,CCp are summarized in the Appendix (for the analogous timings information obtained in the context of the adaptive CC($P$;$Q$) calculations for cyclobutadiene, see Ref. Gururangan and Piecuch, 2023).
[…]